A sound sour moving at a constant speed of 249 m/s towards a detector, emits sound at a constant frequency of 5,100 Hz. Take the speed of sound in the air as 340 m/s. Calculate the wavelength of the sound emitted by the source

1
Expert's answer
2022-04-29T12:30:15-0400

The Doppler's effect equation says

"f=f_0\frac{c}{c-v}=5100*\frac{340}{340-249}=19055\:\rm Hz"

The wavelength

"\lambda=\frac{c}{f}=\frac{340}{19055}=0.018\:\rm m"